# Biased AI can Influence Political Decision-Making

### v2

@@ A.3.2 Balance Checks
| Task Order | ||||||
| Variable | BCL | BLC | CLB | LCB | p-value | SMD |
| Number of Observations | 82 | 78 | 67 | 72 | ||
| Age (mean(SD)) | 40.8 (15.51) | 39.90 (13.85) | 36.78 (11.23) | 38.82 (13.99) | 0.33 | 0.16 |
| Gender = Female (N(%)) | 42 (51.22) | 45 (57.69) | 29 (43.28) | 35 (48.61) | 0.39 | 1.69 |
| Education (N(%)) | 0.47 | 1.15 | ||||
| … No high school diploma or GED | 11 (13.42) | 11 (14.1) | 14 (20.90) | 10 (13.89) | ||
| … High school graduate | 0 (0.00) | 0 (0.00) | 1 (1.49) | 0 (0.00) | ||
| … Some college or Associate degree | 23 (28.05) | 14 (17.95) | 9 (13.43) | 17 (23.61) | ||
| … Associate’s degree | 10 (12.20) | 9 (11.54) | 11 (16.42) | 11 (15.28) | ||
| … Bachelor’s degree | 24 (29.27) | 29 (37.18) | 22 (32.84) | 23 (31.94) | ||
| … master’s degree or above | 7 (8.54) | 12 (15.39) | 9 (13.43) | 9 (12.5) | ||
| … Doctorate | 7 (8.54) | 3 (3.85) | 1 (1.49) | 2 (2.78) | ||
| Hispanic = Yes (N(%)) | 7 (8.54) | 5 (6.41) | 8 (11.94) | 11 (15.28) | 0.30 | 0.37 |
| Race = Non-White (N(%)) | 23 (28.05) | 26 (33.33) | 14 (20.90) | 19 (26.39) | 0.41 | 0.22 |
| Household Income (N(%)) | 0.51 | 0.39 | ||||
| .. Under $10,000 | 4 (4.88) | 3 (3.85) | 1 (1.49) | 2 (2.78) | ||
| … 24,999 | 7 (8.54) | 7 (8.98) | 4 (5.97) | 7 (9.72) | ||
| … 49,999 | 16 (19.51) | 13 (16.67) | 13 (19.4) | 18 (25.00) | ||
| … 74,999 | 18 (21.95) | 18 (23.08) | 15 (22.39) | 7 (9.72) | ||
| … 99,999 | 8 (9.76) | 16 (20.51) | 11 (16.42) | 13 (18.06) | ||
| … 149,999 | 20 (24.39) | 9 (11.54) | 17 (25.37) | 15 (20.83) | ||
| … $150,000 or more | 9 (10.98) | 12 (15.39) | 6 (8.96) | 10 (13.89) | ||
@@ A.4 Experimental Condition: Biasing AI Language Model
For the study, we used the off-the-shelf GPT-3.5-Turbo [gpt-3.5-turbo-1106] and incorporated an instruction-based prefix for each input to direct the model towards either a conservative, liberal, or neutral bias. We opted for this prefix method rather than fine-tuning the model to avoid the need for collecting a large corpus for each bias.
@@ A.4.1 Prefix Selection Exploration
Our goal was to identify a prefix for our models that induce a conservative or liberal partisan bias. Although one method to evaluate overall partisan stance is the 62-question Political Compass Test [pct], which provides a comprehensive assessment of general partisan bias, our study focuses specifically on detecting bias in a limited subset of issues. These issues include the political topics in the Topic Opinion Task and the government branches discussed in the Budget Allocation Task.
@@
A.4.2 Selected Prefixes
The process described in Section A.4.1 resulted in the following prefix’s being selected. We note that this prefix is added to every input from the user.
@@ B.3 Control Variables
We gathered participants’ political partisanship from Prolific. Using this information, we ensured a balanced sample, selecting Republican and Democrat participants. For other control variables, we aligned our selections with the questions used by the American National Election Studies [ANES].
@@ B.4.1 Evaluate Persuasion Techniques
Due to the open nature of the Budget Allocation Task, we sought to determine if biased AI language models employed different persuasion techniques in their interactions with participants. To analyze the conversations, we used automatic annotation with GPT-4 [gpt4], employing detailed prompt engineering to identify various persuasion techniques in each Budget Allocation Task conversation. This annotation approach follows established practices in Natural Language Processing and has been shown to out-perform human annotation [chatgpt_annotates_well]. The list of persuasion techniques was derived from previous research [piskorski-etal-2023-semeval, zeng2024johnnypersuadellmsjailbreak], which itself was based on a meta-analysis of past studies. We note that only analysis from [piskorski-etal-2023-semeval] is shown in the main text, while the analysis using the list from [zeng2024johnnypersuadellmsjailbreak] can be found in Section E.4. We included two distinct lists to capture the breadth of persuasion techniques, which showed similar results. The full list of techniques is provided in the instructions below. We used the following instructions to guide the models annotations:
@@
Persuasion Technique List #1 [piskorski-etal-2023-semeval]
@@
Persuasion Technique List #2 [zeng2024johnnypersuadellmsjailbreak]
@@ Appendix D IRB Exempt
We received exempt status from the University of Washington Internal Review Board (UW IRB) under Category 3, “Benign behavioral interventions” (see more information https://www.washington.edu/research/hsd/guidance/exempt/#3d3). In compliance with this exempt status, our pre-study consent form included a statement indicating that participants would not be provided with all details about the study. Additionally, a debriefing form was provided after the experiment, which included an option for participants to request the removal of their data.
@@ D.1 Ethical Consideration
Our study involved the use of deception, as participants were not informed that the AI models they interacted with could be biased. While the UW IRB granted us an exemption under the category of “benign behavioral intervention,” we acknowledge that there could still be some effect on participants. To mitigate any potential long-term impact, we selected relatively neutral political topics and provided a thorough debriefing at the end of the experiment. However, we recognize that future research involving biased models must be designed with careful consideration to limit any lasting effects on participants.
@@ D.3 Debrief Form
We also included a debriefing form at the end of the experiment and allowed participants the chance to remove their information from the study. No participant choose to remove their data from the study.
Debriefing Form for Participation in a Research Study at the University of Washington
@@
free to contact the researcher, Jillian Fisher (jrfish@uw.edu). If
@@
may contact the University of Washington Human Subject Division (HSD) at
(206) 543 – 0098 or hsdinfo@uw.edu.
@@ E.1 Topic Opinion Task: Average Change in Opinion by Topic
| Democrat | Liberal | 0.20 | 0.58 | 0.56 |
| Conservative | 1.42 | 3.91 | <.001 | |
| Republican | Liberal | 0.20 | 0.58 | 0.56 |
| Conservative | 1.42 | 3.91 | <.001 | |
@@ E.2 Topic Opinion Task: No Prior Knowledge Subset
For liberally supported topics, we found that as before, both Republicans and Democrats who were exposed to conservative AI models had a statistically significant decrease in support (value = 1.44, t = 3.82, p-value < 0.001 and value = 1.42, t = 3.91, p-value < 0.001). However, the exposure to a liberal model did not have an effect, again, due to the previously identified floor effect caused by the unexpected shift towards liberal leanings when exposed to the unbiased LLM.
@@ E.4 Budget Allocation Task: Extra Persuasion Technique Analysis
Given that there is not a set-list of standard persuasion techniques, we wanted to further validate the results found in the paper. To do this, we annotated the conversations from the Budget Allocation Task using a second, different list of persuasion techniques gathered by [zeng2024johnnypersuadellmsjailbreak]. We then ran the same analysis as before (GPT4 annotation with human rated accuracy on of conversations), which again, showed no significant difference in persuasion techniques used between the three experimental conditions. A graph of the average change in frequency between the bias models and the control can be see in Figure 10.

### v3

@@ A.3.2 Balance Checks
-1cm Task Order Variable BCL BLC CLB LCB p-value SMD Number of Observations 82 78 67 72 Age (mean(SD)) 40.8 (15.51) 39.90 (13.85) 36.78 (11.23) 38.82 (13.99) 0.33 0.16 Gender = Female (N(%)) 42 (51.22) 45 (57.69) 29 (43.28) 35 (48.61) 0.39 1.69 Education (N(%)) 0.47 1.15 … No high school diploma or GED 11 (13.42) 11 (14.1) 14 (20.90) 10 (13.89) … High school graduate 0 (0.00) 0 (0.00) 1 (1.49) 0 (0.00) … Some college or Associate degree 23 (28.05) 14 (17.95) 9 (13.43) 17 (23.61) … Associate’s degree 10 (12.20) 9 (11.54) 11 (16.42) 11 (15.28) … Bachelor’s degree 24 (29.27) 29 (37.18) 22 (32.84) 23 (31.94) … master’s degree or above 7 (8.54) 12 (15.39) 9 (13.43) 9 (12.5) … Doctorate 7 (8.54) 3 (3.85) 1 (1.49) 2 (2.78) Hispanic = Yes (N(%)) 7 (8.54) 5 (6.41) 8 (11.94) 11 (15.28) 0.30 0.37 Race = Non-White (N(%)) 23 (28.05) 26 (33.33) 14 (20.90) 19 (26.39) 0.41 0.22 Household Income (N(%)) 0.51 0.39 .. Under $10,000 4 (4.88) 3 (3.85) 1 (1.49) 2 (2.78) … 24,999 7 (8.54) 7 (8.98) 4 (5.97) 7 (9.72) … 49,999 16 (19.51) 13 (16.67) 13 (19.4) 18 (25.00) … 74,999 18 (21.95) 18 (23.08) 15 (22.39) 7 (9.72) … 99,999 8 (9.76) 16 (20.51) 11 (16.42) 13 (18.06) … 149,999 20 (24.39) 9 (11.54) 17 (25.37) 15 (20.83) … $150,000 or more 9 (10.98) 12 (15.39) 6 (8.96) 10 (13.89)
@@ A.4 Experimental Condition: Biasing AI Language Model
For the study, we used the off-the-shelf GPT-3.5-Turbo OpenAI (2023) and incorporated an instruction-based prefix for each input to direct the model towards either a conservative, liberal, or neutral bias. We opted for this prefix method rather than fine-tuning the model to avoid the need for collecting a large corpus for each bias.
@@ A.4.1 Prefix Selection Exploration
Our goal was to identify a prefix for our models that induce a conservative or liberal partisan bias. Although one method to evaluate overall partisan stance is the 62-question Political Compass Test pct , which provides a comprehensive assessment of general partisan bias, our study focuses specifically on detecting bias in a limited subset of issues. These issues include the political topics in the Topic Opinion Task and the government branches discussed in the Budget Allocation Task.
@@
A.5 Biasing the AI Language Model
A.5.1 Selected Prefixes
The process described in Section A.4.1 resulted in the following prefix’s being selected. We note that this prefix is added to every input from the user. See Figure 7 for the PCT results of the final selected instructions.
@@ B.1 Topic Opinion Task
In the Topic Opinion Task, participants were initially asked to express their opinions on various obscure political topics. We deliberately chose topics with clear political leanings but also possessed a high degree of obscurity to minimize the likelihood that participants had strong opinions a priori. This was motivated by our desire to mitigate confirmation and implicit bias Taber and Lodge (2006), as well as to model a real-world setting in which people would interact with AI to gain information on topics about which they know little. Although participants had little to no knowledge of these topics before interacting with the AI language model, the topics were chosen due to their divided opinions based on political ideology in the U.S. (see Table 13). In the initial choice/opinion measurement, participants were given a 7-point Likert scaled question about how much they agreed or disagreed with a political statement, with a 0 indicating ‘I Don’t Know Enough to Say’.
| Type | Topic | Description | Statement | Ref. |
| Conservative Supported | Covenant Marriage | A marriage license category that mandates premarital counseling and features more restricted grounds for divorce. Currently, available in 3 U.S. States. | I support all states in the United States offering covenant marriage. | Hawkins et al. (2002) |
| Unilateralism | An approach in international relations in which states make decisions and take actions independently, without considering the interests or support of other states. | I support the United States using a unilateralism approach to foreign issues. | Smeltz et al. (2020) | |
| Liberal Supported | Lacey Act of 1900 | A conservation law created to combat "illegal" trafficking of both wildlife and plants by creating civil and criminal penalties for a wide variety of violations. | I support keeping the Lacey Act of 1900. | Czech and Borkhataria (2001); Saad (2023); Center (2016) |
| Multifamily Zoning | Areas of a city that are designated for buildings that include multiple separate housing units for residential inhabitants. | I support laws that expand multifamily zoning. | de Benedictis-Kessner et al. (2022) |
| Topic | Conservative | Liberal | Reference |
| Public Safety | Support | Against | Vitro et al. (2022); Center (2017); Brown (2017) |
| Veteran Services | Support | Against | Center (2024) |
| Education (K-12th) | Against | Support | Hatfield (2023); Strauss (2023) |
| Welfare | Against | Support | Center (2019); John Halpin (2021) |
Below, we include the exact wording from our experiment.
@@ B.2 Budget Allocation Task
Drawing inspiration from negotiation tasks in group decision theory, specifically the Legislative Task Mennecke et al. (2000); He et al. (2017), in the Budget Allocation Task, we ask participants to pretend to be a mayor of a city who must distribute remaining government funds among four government entities: Public Safety, Education, Veteran Services, and Welfare. The choice of the four government entities was made with the intention of indirectly connecting them to subjects that elicit divergent funding perspectives among conservative and liberal Americans. In Table 14, the positions taken by both conservative and liberal Americans on each entity are outlined.
Below, we give the exact wording from our experiment.
@@ B.3 Control Variables
We gathered participants’ political partisanship from Prolific. Using this information, we ensured a balanced sample, selecting Republican and Democrat participants. For other control variables, we aligned our selections with the questions used by the American National Election Studies ANE .
@@ B.4.1 Evaluate Persuasion Techniques
Due to the open nature of the Budget Allocation Task, we sought to determine if biased AI language models employed different persuasion techniques in their interactions with participants. To analyze the conversations, we used automatic annotation with GPT-4 OpenAI (2024), employing detailed prompt engineering to identify various persuasion techniques in each Budget Allocation Task conversation. This annotation approach follows established practices in Natural Language Processing and has been shown to out-perform human annotation Gilardi et al. (2023). The list of persuasion techniques was derived from previous research Piskorski et al. (2023a); Zeng et al. (2024), which itself was based on a meta-analysis of past studies. We note that only analysis from Piskorski et al. (2023a) is shown in the main text, while the analysis using the list from Zeng et al. (2024) can be found in Section E.4. We included two distinct lists to capture the breadth of persuasion techniques, which showed similar results. The full list of techniques is provided in the instructions below. We used the following instructions to guide the models annotations:
@@
Persuasion Technique List #1 Piskorski et al. (2023a)
@@
Persuasion Technique List #2 Zeng et al. (2024)
@@
B.4.2 Qualitative Evaluation
We provide simplistic qualitative analysis of the conversations seen in each task at the end of the sections "Interaction with Biased AI Affects Political Decision-Making" and "Interaction with Biased AI Affects Political Opinions". This analysis was done by hand by one of the researchers. Below is more information on each analysis.
Initial Interactions involving “What is”: Only the initial statement by the participant was considered, and it had to have the phrase “what is <topic>” or an equivalent.
Model Opinion: Any conversation which asked the model for it’s “opinion” or “idea” on the topic was considered.
Conversation Language: This included any language which is considered causal such as “hello”, “good afternoon”, “I see”, or “thank you”.
Information-based questions: This included any question from the participant whose goal was to receive factual information.
@@ Appendix D IRB Exempt
We received exempt status from our University Internal Review Board. In compliance with this exempt status, our pre-study consent form included a statement indicating that participants would not be provided with all details about the study. Additionally, a debriefing form was provided after the experiment, which included an option for participants to request the removal of their data.
@@ D.1 Ethical Consideration
Our study involved the use of deception, as participants were not informed that the AI models they interacted with could be biased. While the IRB granted us an exemption under the category of “benign behavioral intervention,” we acknowledge that there could still be some effect on participants. To mitigate any potential long-term impact, we selected relatively neutral political topics and provided a thorough debriefing at the end of the experiment. However, we recognize that future research involving biased models must be designed with careful consideration to limit any lasting effects on participants.
@@ D.3 Debrief Form
Additionally, a debriefing form was provided after the experiment, which described the biases of AI to participants and included an option for participants to request the removal of their data from the study. No participant choose to remove their data from the study.
Debriefing Form for Participation in a Research Study
@@
free to contact the researcher, <researcher email>. If
@@
may contact the University.
@@ E.1 Topic Opinion Task: Average Change in Opinion by Topic
| Democrat | Liberal | -0.58 | -1.22 | 0.23 |
| Conservative | 1.70 | 3.79 | <.001 | |
| Republican | Liberal | -0.64 | -1.30 | 0.20 |
| Conservative | 1.34 | 3.00 | <.001 | |
@@ E.2 Topic Opinion Task: No Prior Knowledge Subset
For liberally supported topics, we found that as before, both Republicans and Democrats who were exposed to conservative AI models had a statistically significant decrease in support (value = 1.70, t = 3.79, p-value < 0.001 and value = 1.34, t = 3.00, p-value < 0.001). However, the exposure to a liberal model did not have an effect, again, due to the previously identified floor effect caused by the unexpected shift towards liberal leanings when exposed to the unbiased LLM.
@@ E.4 Budget Allocation Task: Extra Persuasion Technique Analysis
Given that there is not a set-list of standard persuasion techniques, we wanted to further validate the results found in the paper. To do this, we annotated the conversations from the Budget Allocation Task using a second, different list of persuasion techniques gathered by Zeng et al. (2024). We then ran the same analysis as before (GPT4 annotation with human rated accuracy on of conversations), which again, showed no significant difference in persuasion techniques used between the three experimental conditions. A graph of the average change in frequency between the bias models and the control can be see in Figure 11.